OpenStudy (anonymous):

how does 3sqrt8 = 2 ?

OpenStudy (anonymous):

\[2\times 2\times 2=2^3=8\]

OpenStudy (anonymous):

its the cube root

OpenStudy (anonymous):

wouldn't sqrt of 8 have to = 2/3?

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

\[\sqrt{8}=\sqrt{4\times 2}=2\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

\[\sqrt[3]{8}=\sqrt[3]{2^3}=2\], take note that \[\sqrt[3]{x^3}=x\] for any number x. Basically, the cube root "undoes" the cube exponent. So if you cube 2, then take the cube root, you'll get back 2 as an answer.

OpenStudy (anonymous):

thank u